A frustum of a square pyramid has an upper base edge of 6 cm and a lower base edge of 4 cm if the distance between these 2 bases is 10 cm. Find the total surface area of the frustum.


1
Expert's answer
2021-08-19T08:00:51-0400


let small height, h = xcm

let big height, H=(10+x)

"\frac{10+x}{x}=\frac{6}{4}"

"40+4x=6x"

"2x=40"

"x=h=20"

H"=10+20=30cm"

for the small pyramid, height of one face "=\sqrt{20^2+2^2}=20.1cm"

"=\sqrt{20^2+2^2}=20.1cm"

"area\space of \space face=1/2\times 4\times 20.1=40.2cm^2"

area of the four faces "=40.2\times 4=160.8cm^2"

for the big pyramid, height of one face="\sqrt{30^2+3^2}=30.15cm"

"\sqrt{30^2+3^2}=30.15cm"

area of the face "=1/2\times 6\times 30.15=90.45cm^2"

area of four faces"=90.45\times 4=361.8cm^2"

area of sides of frustum=361.8-160.8=201

assuming the frustum is closed both sides area of upper base "=6\times 6=36cm^2"

area of lower base"=4\times 4=16cm^2"

total area of frustum="201+36+16=253cm^2"